OpenStudy (anonymous):

How many photons are produced in a laser pulse of 0.655 J at 607 nm?

OpenStudy (schrodingers_cat):

E= ch/(wavelength) E= Energy per photon So, E= (6.626x10^-34)(3.00x10^8)/(607x10^-9)= 3.27 x10^-19 j Then, .655/(3.27x10^-19) = 2.00x10^18 photons Hope this helps Might wanna run through it yourself just to double check numbers :)